OpenStudy (anonymous):

There is a 70% chance of heavy fog and a 30% chance of light drizzle. A softball game has an 80% chance of being cancelled for a light drizzle and a 30% chance of being cancelled for a heavy for when there is no drizzle. Find the probability the game will be cancelled.

OpenStudy (anonymous):

Prob of heavy fog , H = 0.7 Prob of light drizzle, L = 0.3 Prob of game cancelled during H, X= 0.3 Prob of game cancelled during L, Y = 0.8 Prob of game cancelled = XH + YL = 0.21 + 0.24 = 0.45